Can you please help me solve for "x" : 2x/3=2
Answer by rothauserc(4718)

solve for "x" : 2x/3=2
multiply both sides of = by 3
2x = 6
x = 3
check the answer
2*3 / 3 = 2
6 / 3 = 2
2 = 2